A man moves on a straight horizontal road with a block in his hand . If he covers a distance of 40m with an acceleration of 0.5ms^-2. find the work done during the motion

Grade:11

3 Answers

Shailendra Kumar Sharma
6 years ago
assume the mass of block be m 
so work done can be calculated W =change in kinetic energy =K2-K1
the velocity after covering 40m 
V2=u2+2aS
So V2 =40
Now kinetic energy will be (1/2)mV2 =20m
So the work done will be 20mJ
Gurmeet
6 years ago
Let suppose a block of mass m so work energy theoryWork done=change in kinetic energy =K2-K1the velocity attined after covering 40m V2=u2+2aSSo V2 =40Now change in kinetic energy will be (1/2)mV2 =20mSo the work done will be 20mJ using above equations
Sonam Tyagi
5 years ago
Given that 
a = 0.5 ms^2
m = 2 kg
s = 40 m 
Now to find v^2 we use
v^2= u^2 + 2as
v= √40
Since by work energy therom
Work done=∆ K.E.
W= 40J